please help me solve this.
2/3 (a-3)= -1/3
please show me how you do it..

Answer by LilSkittleMd(119)

2/3 (a-3)= -1/3
Distribute
2/3a-2=-1/3
Add 2 to both sides of the equation
2/3a=5/3
Divide each side by 2/3
a=5/2